# Few-Class Arena: A Benchmark for Efficient Selection of Vision Models and Dataset Difficulty Measurement

### v1

@@ 4.2 Results on FC-Sub
The preceding observation for FC-Full 4.1 seems to contradict the common belief that, the fewer the classes, the higher is the accuracy that a model can achieve. Conversely, the FC-Sub results do align with this belief. We argue that a full-class model needs to accommodate many parameters to learn features that will enable high performance across all classes in a many-class, full dataset. With the same parameters, however, a sub-class model can adapt to finer and more discriminative features that improve its performance when the number of target classes are much smaller.
@@ 4.3 Results on FC-Sim
In our experiments, we leverage the general visual feature extraction ability of CLIP (image + text) [52] and DINOv2 (image) [53] by self-supervised learning. Specifically, a pair of images are fed into its latent space from which the the cosine score is calculated and normalized to 00 to $1$. Note that we only use the Image Encoder in CLIP.
Comparing Accuracy and Similarity To evaluate SimSS, we compute the Pearson correlation coefficient (PCC) ($r$) between model accuracy and SimSS. Results in Figure 5 (a) (b) show that SimSS is poorly correlated with DCN-Full ($r=0.18$ and $r=0.26$ for CLIP and DINOv2) due to the large variance shown in Section 4.1. In contrast, SimSS is highly correlated with DCN-Sub (shown in blue squares), with $r=0.90$ and $r=0.88$ using CLIP (dashed) and DINOv2 (solid), respectively. The high PCC [62, 63] demonstrates that SimSS is a reliable metric to estimate few-class dataset difficulty, and this can help predict the empirical upper-bound accuracy of a model in the Few-Class Regime. Comparison between SimSS and all models can be found in the Appendix. Such a high correlation suggests this offers a reliable scaling relationship to estimate model accuracy by similarity for other values of $N_{CL}$ without an exhaustive search. Due to the dataset specificity of the dataset difficulty property, this score is computed once and used for all times the same dataset is used. We have made available difficulty scores for many datasets at the Few-Class Arena site.
@@ 5 Conclusion
We have proposed Few-Class Arena and a dataset difficulty measurement, which together form a benchmark tool to compare and select efficient models in the Few-Class Regime. Extensive experiments and analyses over 1500 experiments with 10 models on 10 datasets have helped identify new behavior that is specific to the Few-Class Regime as compared to for many-classes. One finding reveals a new $N_{CL}$-scaling law whereby dataset difficulty must be taken into consideration for accuracy prediction. Such a benchmark will be valuable to the community by providing both researchers and practitioners with a unified framework for future research and real applications.
Limitations and Future Work. We note that the convergence of sub-models is contingent on various factors in a training scheduler, such as learning rate. A careful tuning of training procedure may increase a model’s performance, but it shouldn’t change the classification difficulty number drastically since this represents a dataset’s intrinsic difficulty property. The current difficulty benchmark supports image similarity while in the future it can be expanded to other difficulty measurements [25, 64]. As CLIP and DINOv2 are trained toward general visual features, it is unclear if they will be appropriate for other types of images such as sketches without textures in Quickdraw [65] . For this reason, a universal similarity foundation model would be appealing that applies to any image type. In summary, Few-Class Arena identifies a promising new path to achieve efficiencies that are focused on the important and practical Few-Class Regime, establishing this as a baseline for future work.
@@ Appendix A Appendix
A.1 Full Models on ImageNet
In practice, ImageNet serves as a common benchmark for vision neural networks. We list the details of 10 pre-trained models from MMPreTrain [55] in terms of Top-1 Accuracy and scale (#$\#Params$) in Fig. 6.
@@ A.1 Full Models on ImageNet
Conformer [66] ConvNeXt [67]
EfficientFormer [68] EfficientNet [69]
RepVGG [70] ResNet [71]
ShuffleNet [72] Swin [73]
VGG [74] ViT [75]
@@
A.2 Extended Many-Class Full Dataset Trained Benchmark Results
A complete ranking of 10 models in 10 datasets is depicted in Fig. 7. Observe that the 10 models’ rankings differ dramatically among 10 different datasets where each line changes from ImageNet1K (IN1K) to other datasets. This poses some questions whether rankings in existing benchmarks can be a reliable indicator for a practitioner to select an efficient neural network, especially when the deployed environment changes from application to application. A major variable in this process is the reduced number of classes from benchmark datasets to deployed environments in the Few-Class Regime. As such, our tool is developed to facilitate research in the Few-Class Regime.
Figure 7: Extended Details of Fig. 2 (b) in the main paper. Full Ranking of 10 models across 10 datasets by Top-1 acc.
A.3 Datasets
@@
Caltech 101 CT101 [76] data.caltech.edu/records/mzrjq-6wc02 tools/ncls/datasets/caltech101.py
Caltech 256 CT256 [77] data.caltech.edu/records/nyy15-4j048 tools/ncls/datasets/caltech256.py
CIFAR-100 CF100 [78] cs.toronto.edu/ kriz/cifar.html tools/ncls/datasets/cifar100.py
@@
Caltech-UCSD CB200 [79] vision.caltech.edu/visipedia/CUB-200-2011.html tools/ncls/datasets/cub200.py
@@
Food 101 FD101 [80] vision.ee.ethz.ch/datasets_extra/food-101/ tools/ncls/datasets/food101.py
@@
German Traffic Sign GT43 [81] benchmark.ini.rub.de/ tools/ncls/datasets/gtsrb43.py
ImageNet IN1K [1] image-net.org/challenges/LSVRC/2012/index.php *
Indoor Scene Recognition ID67 [82] web.mit.edu/torralba/www/indoor.html tools/ncls/datasets/indoor67.py
Quickdraw QD345 [65] github.com/googlecreativelab/quickdraw-dataset tools/ncls/datasets/quickdraw345.py
@@
Describable Textures Dataset TT47 [83] robots.ox.ac.uk/ vgg/data/dtd/index.html tools/ncls/datasets/textures47.py
Table 3: Dataset information. * Note that ImageNet dataset format is used as the reference for other datasets. Therefore Path in FCA for ImageNet is not required.
@@
where a .txt file stores a pair of image id and and class number in each row in the following format
@@
A.4 Model Training Details
@@
ResNet50 RN50 [71] SGD 0.1 1e-4 momentum=0.9
VGG16 VGG16 [84] SGD 0.01 1e-4 momentum=0.9
ConvNeXt V2 CNv2 [85] AdamW 2.5e-3 0.05 eps=1e-8
@@
Inception V3 INCv3 [86] SGD 0.1 1e-4 momentum=0.9
EfficientNet V2 EFv2 [4] SGD 4e-3 0.1 momentum=0.9
@@
ShuffleNet V2 SNv2 [9] SGD 0.5 0.9 momentum=4e-5
MobileNet V3 MNv3 [7] RMSprop 6.4e-4 1e-5 alpha=0.9
@@
Vision Transformer ViTb [75] AdamW 3e-3 0.3 -
@@
Swin Transformer V2 SWv2b [87] AdamW 1e-4 0.05 eps=1e-8
@@
MobileViT MViTs [10] SGD 0.1 1e-4 momentum=0.9
@@
Table 5: Model Training information. LR: Learning rate. SGD: Stochastic gradient descent. AdamW: Adam with weight decay. RMSprop: Root mean square propagation.
A.5 Extended Few-Class Full Dataset Trained Benchmark (FC-Full) Results
@@
A.6 Extended Few-Class Similarity Benchmark (FC-Sim) Details
We present the extended mathematical details of 3.6 Few-Class Similarity Benchmark (FC-Sim) in the main paper.
The basic similarity formulation is adopted from [51]. Notations are defined as follows:
@@ A.6 Extended Few-Class Similarity Benchmark (FC-Sim) Details
$S_{\alpha}^{(C)}=\frac{1}{|P^{(C)}|}\sum_{i,j\in C;i\neq j}\cos(\textbf{Z}_{i},\textbf{Z}_{j}).$ (5)
@@
$S_{\alpha}^{(D)}=\frac{1}{|L|}\sum_{l\in L}S_{\alpha}^{(C_{l})}=\frac{1}{|L|\times|P^{(C_{l})}|}\sum_{l\in L}\sum_{i,j\in C_{l};i\neq j}\cos(\textbf{Z}_{i},\textbf{Z}_{j}),$ (7)
$S_{\beta}^{(D)}=\frac{1}{|P^{(D)}|}\sum_{a,b\in L;a\neq b}S_{\beta}^{(C_{a},C_{b})}=\frac{1}{|P^{(D)}|\times|P^{(C_{1},C_{2})}|}\sum_{a,b\in L;a\neq b}\sum_{i\in C_{1},j\in C_{2}}\cos(\textbf{Z}_{i},\textbf{Z}_{j}).$ (8)
@@
To compensate the aforementioned drawback, we adopt the Silhouette Score (SS) (also called Silhouette Coefficient in the literature) [34, 61]:
@@
where $SS(i)$ is the Silhouette Score of the data point $i$, $a(i)$ is the average dissimilarity between $i$ and other instances in the same class, and $b(i)$ is the average dissimilarity between $i$ and other data points in the closest different class. Intuitively, this metric summarizes the quality of clusters jointly by the degree of instances of the same class and distinct clusters, normalized by the longest distance of $a(i)$ and $b(i)$. By this definition, we can see that $-1\leq SS(i)\leq 1$ where $-1$ indicates a dataset is poorly clustered (data points with different classes are scattered around) while $1$ represents a well-clustered dataset.
Euclidean Distance is commonly used to measure two data points’ differences; in contrast, we incorporate the inverse of similarity (dissimilarity) as data point’s difference into the existing Silhouette Score. Observe that the above Intra-Class Similarity $S_{\alpha}^{(C)}$ already represents the tightness of the class $(C)$, therefore $a(i)$ can be replaced with the inverse of Intra-Class Similarity $a(i)=-S_{\alpha}(i)$. For the second term $b(i)$, we adopt the previously defined Inter-Class Similarity $S_{\beta}^{(C_{1},C_{2})}$ and introduce a new similarity score as follows:
@@
$S^{\prime(C)}_{\beta}=\frac{1}{|P^{(C,\hat{C})}|}\sum_{i\in C,j\in\hat{C}}\cos(\textbf{Z}_{i},\textbf{Z}_{j}),$ (10)
where $\hat{C}$ is the set of the nearest class to $C$ ($\hat{C}\neq C$).
Consequently, the dataset-level Nearest Inter-Class Similarity $S^{\prime(D)}_{\beta}$ is expressed as:
$S^{\prime(D)}_{\beta}=\frac{1}{|L|}\sum_{l\in L}S^{\prime(C_{l},\hat{C_{l}})}_{\beta}=\frac{1}{|L|\times|P^{(C_{l},\hat{C_{l}})}|}\sum_{l\in L}\sum_{i\in C_{l},j\in\hat{C_{l}}}\cos(\textbf{Z}_{i},\textbf{Z}_{j}).$ (11)
@@
A.7 Extended Few-Class Similarity Benchmark (FC-Sim) Results

### v2

@@ 4.2 Results on FC-Sub
The preceding observation for FC-Full 4.1 seems to contradict the common belief that, the fewer the classes, the higher the accuracy a model can achieve. Conversely, the FC-Sub results do align with this belief. We argue that a full-class model needs to accommodate many parameters to learn features that will enable high performance across all classes in a many-class, full dataset. With the same parameters, however, a sub-class model can adapt to finer and more discriminative features that improve its performance when the number of target classes are much smaller.
@@ 4.3 Results on FC-Sim
In our experiments, we leverage the general visual feature extraction ability of CLIP (image + text) (Radford et al., 2021) and DINOv2 (image) (Oquab et al., 2023) by self-supervised learning. Specifically, a pair of images are fed into its latent space from which the the cosine score is calculated and normalized to 00 to $1$. Note that we only use the Image Encoder in CLIP.
Comparing Accuracy and Similarity To evaluate SimSS, we compute the Pearson correlation coefficient (PCC) ($r$) between model accuracy and SimSS. Results in Figure 5 (a) (b) show that SimSS is poorly correlated with DCN-Full ($r=0.18$ and $r=0.26$ for CLIP and DINOv2) due to the large variance shown in Section 4.1 , as well as the general features learned on a full dataset, those of which can be extraneous to a much smaller set of target sub-class features. In contrast, SimSS is highly correlated with DCN-Sub (shown in blue squares), with $r=0.90$ and $r=0.88$ using CLIP (dashed) and DINOv2 (solid), respectively. We attribute the advantages of DCN-Sub to its focus on the minimal features tailored to the target sub-classes, while maintaining the same number of parameters as the DCN-Full architecture. The high PCC (Wicklin, 2024; Schober et al., 2018) demonstrates that SimSS is a reliable metric to estimate few-class dataset difficulty, and this can help predict the empirical upper-bound accuracy of a model in the Few-Class Regime. Comparison between SimSS and all models can be found in the Appendix. Such a high correlation suggests this offers a reliable scaling relationship to estimate model accuracy by similarity for other values of $N_{CL}$ without an exhaustive search. Due to the dataset specificity of the dataset difficulty property, this score is computed once and used for all times the same dataset is used. We have made available difficulty scores for many datasets at the Few-Class Arena site.
@@
4.4 Comparison with Fine-Tuned Models
Fine-tuning a model through transfer learning from a pre-trained model has become as a common practice in many real-world scenarios. We perform experiments on CNNs (ResNet18, 50) and Transformer architectures (MobileVit-small (Mehta & Rastegari, 2021), ViT-Base) summarized in Table 2. To fine-tune (FT) the ResNet18, 50 and MobileVit-small models for $N_{CL}\in\{2,4\}$, we first train their full models $N_{CL}=\{100\}$ on CIFAR100 for 100 epochs, and subsequently fine-tune their weights for the target $N_{CL}\in\{2,4\}$ for another 20 epochs. For ViT, we fine-tune a ViT-B model initialized with weights from the CLIP pre-trained backbone. A linear layer is added on top, and the model is trained for 10 epochs. This setup is indicated by the star symbol (*). We conclude that the fine-tuned models exhibit patterns and trends consistent with the observations presented in Fig. 1. Note our focus of this work is to leverage the proposed difficult measurement method, FC-Sim, to efficiently estimate the achievable model accuracy, thereby assisting in model selection in the Few-Class Regime. Sub-models can offer insights into the minimal visual features required for a specific real-world scenario as they are trained exclusively on the target classes. In contrast, weights pre-trained on large full datasets – whether through fully supervised manner or self-supervised – may include extraneous features that are irrelevant to the target classes. We therefore prioritize sub-model study in this work.
MT $N_{CL}$ ResNet18 ResNet50 MViT-S ViT-B
F 100 76.11 73.71 73.83 32.54
F 4 75.10 72.20 72.35 36.15
FT 4 87.60 90.55 90.00   91.16*
S 4 90.65 90.15 89.45 85.40
F 2 75.00 71.30 71.80 40.80
FT 2 87.90 93.70 90.50   95.20*
S 2 96.30 95.30 95.50 95.90
Table 2: Top-1 Accuracies for different configurations on CIFAR100. $N_{CL}\in\{2,4,100\}$, MT: Model Type, F: Full model, S: Sub-model, FT: Fine-tuned model, MViT-S: MobileViT-small, *: fine-tuned from the CLIP pre-trained model. Best scores are highlighted in bold. The gray bar indicates sub-models as the primary focus of this research.
@@ 5 Conclusion
We have proposed Few-Class Arena and a dataset difficulty measurement, which together form a benchmark tool to compare and select efficient models in the Few-Class Regime. Extensive experiments and analyses over 1500 experiments with ten models on ten datasets have helped identify new behaviors specific to the Few-Class Regime as compared to many-classes setting. One finding reveals a new $N_{CL}$-scaling law whereby dataset difficulty must be taken into consideration for accuracy prediction. Such a benchmark will be valuable to the community by providing both researchers and practitioners with a unified framework for future research and real applications.
Limitations and Future Work. The current difficulty benchmark supports image similarity while in the future it can be expanded to other difficulty measurements (Scheidegger et al., 2021). CLIP and DINOv2 are trained toward general visual features, it is unclear if they will be appropriate for other types of images such as sketches without textures in Quickdraw (Ha & Eck, 2017). For this reason, a universal similarity foundation model would be appealing that applies to any image type. Moreover, integrating image similarity with representation similarity (AntixK, 2023; Nguyen et al., 2020; Cao et al., 2024) could further enhance model efficiency, leveraging complementary insights from both approaches. In summary, Few-Class Arena identifies a promising new path for achieving efficiencies focused on the important and practical Few-Class Regime, establishing a baseline for future work.
Acknowledgments
This research has been supported in part by the National Science Foundation (NSF) under Award numbers 2055520 and 2238553. We appreciate the GPU-cluster maintenance support from Rob Dinoff and Hans Woithe for running large-scale experiments, as well as insightful discussions with Zhengyu Wu and Xi Han significantly enhanced the clarity and readability of our plots.
@@ Appendix A Appendix
A.1 Goals
A.2 Benchmark Usage Guideline
Users should prepare the dataset detailed in 3.1 in the “CUSTOM“ format based on the MMPreTrain (Contributors, 2023) documentation. The tool is designed for both practitioners and researchers.
Practitioner: Users select the target $N_{CL}$ and then execute the FC-Sim on the custom dataset. FC-Sim calculates the image similarity score which can be used to index a narrow range of potential target models, given the deployment accuracy requirement (e.g. accuracy). We provide the index table covering ten models on ten datasets for the Few-Class Regime.
Researcher: Users specify the configurations for FC-Full, FC-Sub and FC-Sim, as well as a list of $N_{CL}$. In the next step, users can execute the scripts running on all benchmarks on the custom dataset. The results for all benchmarks are then used for further analysis.
A.3 Extended Related Work
Few-Shot Learning. There has been a large body of research on Few-Short Learning (FSL) (Song et al., 2023; Wang et al., 2020; Hu et al., 2022; Sung et al., 2018). However, the fundamental research questions differ from ours in the Few-Class Regime. The FSL framework aims to address the problem of data scarsity with the goal for a model to leverage the representations from very few samples (or none, in the case of Zero-Shot Learning), or prior knowledge that can generalize effectively to other tasks or domains.
Self-Supervised Learning. To leverage the knowledge from unlabled data, Self-Supervised Learning (SSL) has emerged as an effective learning framework to learn general vision features (Jaiswal et al., 2020; Jing & Tian, 2020). This includes techniques such as Contrastive Learning, applied to single modalities (Chen et al., 2020a; Chen & He, 2021; He et al., 2020; Chen et al., 2020b; 2021) or multiple modalities (Radford et al., 2021), as well as mask-and-reconstruct methods (He et al., 2022), among others.
In contrast to the advancements of the aforementioned learning frameworks, Few-Class Arena focuses on the research problem of selecting the most efficient model with minimal features (e.g. model parameters, model multiplies) needed for the target application deployment.
A.4 Full Models on ImageNet
In practice, ImageNet serves as a common benchmark for vision neural networks. We list the details of ten pre-trained models from MMPreTrain (Contributors, 2023) in terms of Top-1 Accuracy and scale (#$\#Params$) in Fig. 6.
@@ A.4 Full Models on ImageNet
Conformer (Peng et al., 2021) ConvNeXt (Liu et al., 2022b)
EfficientFormer (Li et al., 2022b) EfficientNet (Tan & Le, 2019b)
RepVGG (Ding et al., 2021) ResNet (He et al., 2016)
ShuffleNet (Zhang et al., 2017) Swin (Liu et al., 2021)
VGG (Simonyan & Zisserman, 2014a) ViT (Dosovitskiy et al., 2020)
@@
A.5 Extended Many-Class Full Dataset Trained Benchmark Results
A complete ranking of ten models in ten datasets is depicted in Fig. 7. Observe that the ten models’ rankings differ dramatically among ten different datasets where each line changes from ImageNet1K (IN1K) to other datasets. This poses some questions whether rankings in existing benchmarks can be a reliable indicator for a practitioner to select an efficient neural network, especially when the deployed environment changes from application to application. A major variable in this process is the reduced number of classes from benchmark datasets to deployed environments in the Few-Class Regime. As such, our tool is developed to facilitate research in the Few-Class Regime.
Figure 7: Extended Details of Fig. 2 (b) in the main paper. Full Ranking of ten models across ten datasets by Top-1 acc.
A.6 Datasets
@@
Caltech 101 CT101 (Li et al., 2022a) data.caltech.edu/records/mzrjq-6wc02 tools/ncls/datasets/caltech101.py
Caltech 256 CT256 (Griffin et al., 2022) data.caltech.edu/records/nyy15-4j048 tools/ncls/datasets/caltech256.py
CIFAR-100 CF100 (Krizhevsky et al., 2009) cs.toronto.edu/ kriz/cifar.html tools/ncls/datasets/cifar100.py
@@
Caltech-UCSD CB200 (Wah et al., 2011) vision.caltech.edu/visipedia/CUB-200-2011.html tools/ncls/datasets/cub200.py
@@
Food 101 FD101 (Bossard et al., 2014) vision.ee.ethz.ch/datasets_extra/food-101/ tools/ncls/datasets/food101.py
@@
German Traffic Sign GT43 (Stallkamp et al., 2012) benchmark.ini.rub.de/ tools/ncls/datasets/gtsrb43.py
ImageNet IN1K (Deng et al., 2009) image-net.org/challenges/LSVRC/2012/index.php *
Indoor Scene Recognition ID67 (Quattoni & Torralba, 2009) web.mit.edu/torralba/www/indoor.html tools/ncls/datasets/indoor67.py
Quickdraw QD345 (Ha & Eck, 2017) github.com/googlecreativelab/quickdraw-dataset tools/ncls/datasets/quickdraw345.py
@@
Describable Textures Dataset TT47 (Cimpoi et al., 2014) robots.ox.ac.uk/ vgg/data/dtd/index.html tools/ncls/datasets/textures47.py
Table 4: Dataset information. * Note that ImageNet dataset format is used as the reference for other datasets. Therefore, the Path in FCA is not required for ImageNet.
@@
where a .txt file stores a pair of image id and and class number in each row in the following format:
@@
A.7 Model Training Details
@@
ResNet50 RN50 (He et al., 2016) SGD 0.1 1e-4 momentum=0.9
VGG16 VGG16 (Simonyan & Zisserman, 2014b) SGD 0.01 1e-4 momentum=0.9
ConvNeXt V2 CNv2 (Woo et al., 2023) AdamW 2.5e-3 0.05 eps=1e-8
@@
Inception V3 INCv3 (Szegedy et al., 2016) SGD 0.1 1e-4 momentum=0.9
EfficientNet V2 EFv2 (Tan & L., 2021) SGD 4e-3 0.1 momentum=0.9
@@
ShuffleNet V2 SNv2 (Ma et al., 2018) SGD 0.5 0.9 momentum=4e-5
MobileNet V3 MNv3 (Howard et al., 2019) RMSprop 6.4e-4 1e-5 alpha=0.9
@@
Vision Transformer ViTb (Dosovitskiy et al., 2020) AdamW 3e-3 0.3 -
@@
Swin Transformer V2 SWv2b (Liu et al., 2022a) AdamW 1e-4 0.05 eps=1e-8
@@
MobileViT MViTs (Mehta & Rastegari, ) SGD 0.1 1e-4 momentum=0.9
@@
Table 6: Model Training detials. LR: Learning rate. SGD: Stochastic gradient descent. AdamW: Adam with weight decay. RMSprop: Root mean square propagation.
A.8 Extended Few-Class Full Dataset Trained Benchmark (FC-Full) Results
@@
A.9 Extended Few-Class Similarity Benchmark (FC-Sim) Details
We present the extended mathematical details of Section 3.5 Few-Class Similarity Benchmark (FC-Sim) in the main paper.
The basic similarity formulation is adopted from (Cao et al., 2023). Notations are defined as follows:
@@ A.9 Extended Few-Class Similarity Benchmark (FC-Sim) Details
$S_{\alpha}^{(C)}=\frac{1}{|P^{(C)}|}\sum_{i,j\in C;\hskip 2.0pti\neq j}\cos(\textbf{Z}_{i},\textbf{Z}_{j}).$ (5)
@@
$S_{\alpha}^{(D)}=\frac{1}{|L|}\sum_{l\in L}S_{\alpha}^{(C_{l})}=\frac{1}{|L|\times|P^{(C_{l})}|}\sum_{l\in L}\hskip 2.0pt\sum_{i,j\in C_{l};\hskip 2.0pti\neq j}\cos(\textbf{Z}_{i},\textbf{Z}_{j}),$ (7)
$S_{\beta}^{(D)}=\frac{1}{|P^{(D)}|}\sum_{a,b\in L;a\neq b}S_{\beta}^{(C_{a},C_{b})}=\frac{1}{|P^{(D)}|\times|P^{(C_{1},C_{2})}|}\sum_{a,b\in L;\hskip 2.0pta\neq b}\hskip 2.0pt\sum_{i\in C_{1},j\in C_{2}}\cos(\textbf{Z}_{i},\textbf{Z}_{j}).$ (8)
@@
To compensate the aforementioned drawback, we adopt the Silhouette Score (SS) (also called Silhouette Coefficient in the literature) (Rousseeuw, 1987; Shahapure & Nicholas, 2020):
@@
where $SS(i)$ is the Silhouette Score of the data point $i$, $a(i)$ is the average dissimilarity between $i$ and other instances in the same class, and $b(i)$ is the average dissimilarity between $i$ and other data points in the closest different class. Intuitively, this metric summarizes the quality of clusters by jointly considering the distance between instances of the same class and distinct clusters, normalized by the longest distance of $a(i)$ and $b(i)$. By this definition, we can see that $-1\leq SS(i)\leq 1$ where $-1$ indicates a dataset is poorly clustered (data points with different classes are scattered around) while $1$ represents a well-clustered dataset.
Euclidean Distance is commonly used to measure two data points’ differences; in contrast, we incorporate the inverse of similarity (dissimilarity) as data points’ differences into the existing Silhouette Score. Observe that the above Intra-Class Similarity $S_{\alpha}^{(C)}$ already represents the tightness of the class $(C)$, therefore $a(i)$ can be replaced with the inverse of Intra-Class Similarity $a(i)=-S_{\alpha}(i)$. For the second term $b(i)$, we adopt the previously defined Inter-Class Similarity $S_{\beta}^{(C_{1},C_{2})}$ and introduce a new similarity score as follows:
@@
${S^{\prime}}_{\beta}^{(C)}=\frac{1}{|P^{(C,\hat{C})}|}\sum_{i\in C,j\in\hat{C}}\cos(\textbf{Z}_{i},\textbf{Z}_{j}),$ (10)
where $\hat{C}$ is the nearest class to instance $i$ ($\hat{C}\neq C$). To determine $\hat{C}$, we first iterate over all other classes $C^{\prime}$ that are different from $C$. For each $C^{\prime}$, we compute the average similarity between $i$ and all samples in $C^{\prime}$. The class $C^{\prime}$ with the highest average similarity score is then chosen as $\hat{C}$.
Consequently, the dataset-level Nearest Inter-Class Similarity ${S^{\prime}}_{\beta}^{(D)}$ is expressed as:
${S^{\prime}}_{\beta}^{(D)}=\frac{1}{|L|}\sum_{l\in L}{S^{\prime}}_{\beta}^{(C_{l})}=\frac{1}{|L|\times|P^{(C_{l},\hat{C_{l}})}|}\sum_{l\in L}\hskip 2.0pt\sum_{i\in C_{l},j\in\hat{C_{l}}}\cos(\textbf{Z}_{i},\textbf{Z}_{j}).$ (11)
@@
A.10 Extended Few-Class Similarity Benchmark (FC-Sim) Results